# GRU: Gated Recurrent Unit with Spatial Attention for Spatio-Temporal Disaggregation

### v3

@@ 3.1. Data Processing
We first define a 128256 grid and plot the regions at each geographic level. Each region has a corresponding value from the vector input ā , and the number of pixels located within the region ā . Then each pixel in each region takes the value of . Pixels shared by more than one region are assigned to whichever region is smaller to avoid ambiguity. This representation captures the shape of, and spatial relationships between, individual regions and affords the use of convolutional layers to model local and non-local influences. The cost is that we lose information during translation since the raster representation does not represent region boundaries precisely, especially at BLOCK level.
We compute descriptive statistics, including means and standard deviations, for the counts of each dataset at different geographic levels. These statistics are shown in Table 1. For NYC Taxi data, the standard deviations are large demonstrating the significant variability in the count values between regions. For instance, Inwood in upper Manhattan has considerably fewer taxi rides compared to the more central areas of Manhattan. In the case of the NYC Bikeshare, we notice significantly smaller average values at the TRACT and BLOCK levels. This observation can be attributed to the distribution of the bikes, which may vary in frequency across different regions within the city. As a result, the average values of these datasets are lower at the more granular levels. For both datasets, we use the month of June as the test split, the month of May as the validation split, and the rest as training data.
| Data | PUMA | NTA | TRACT | BLOCK |
| NYC Taxi | 1443.01 | 450.94 | 51.17 | 3.86 |
| (1776.03) | (689.16) | (73.32) | (8.03) | |
| NYC Bikeshare | 189.24 | 59.14 | 6.71 | 0.50 |
| (279.28) | (97.86) | (13.09) | (2.51) |
@@ 4. Models
In this section, we describe the GRU model composed of spatial attention layers within a GRU framework. We also describe the containment property that relates different levels of the aggregation hierarchy that can be used for supervision.
Notation We adopt the following notation:
ā¢
@@
4.1. Containment Property
@@
The geographic levels we use exhibit containment: each region at level is wholly contained in a region at level such that regions are hierarchical. This property tends to hold in political contexts. For example, one census tract comprises multiple census blocks. Assuming containment, we can identify a one-to-many map among the regions from any two geographic levels. Let be a binary matrix representing the containment relationship between geographic levels and . A value of 1 at position indicates that region at level is contained in region at level . Figure 2(a) shows an example where three regions from level at a single th timestamp are split into six regions from level . Figure 2(b) presents the corresponding containment map.
4.2. Spatial Attention (SA) Layer
@@
Spatio-temporal transformers have been used in urban prediction tasks (§2). The spatial attention layer in the transformer block can capture complex spatial relationships among different regions to improve performance. In the disaggregation task, spatial attention is calculated across two different geographic levels (Figure 2(c)). A spatial attention head () for the disaggregation task works as follows ā Given one timestamp input from geographic level , we first calculate the key, query, and value metrics:
and have the same dimension as level , while has the dimension of the geographic level . Then we attend the coarse region values to the finer region values by calculating the dot-product spatial attention. Additionally, we use the binary containment maps to force the attention to happen only among the connected regions:
The containment maps represent a constraint that child region values must sum to their parentās region value, ensuring coherent and consistent values across different geographic resolutions. The disaggregation output from a single spatial attention head is then calculated as:
@@
A spatial attention () layer is composed with multiple spatial attention heads , enhancing the model capacity (Figure 3):
The same calculation works with temporal inputs. Given hours of input from geographic level , we reshape the dimension and calculate spatial attention based on Eq (6).
@@
In the typical spatial attention calculationĀ (Vaswani etĀ al., 2017), a residual connection layer is applied after the multi-head attention layer: . In the disaggregation task, the input and output have different dimensions, which cannot be added. Therefore, residual connection is not applicable. In our experiment, each layer uses eight .
4.3. GRU
@@
The spatial attention layer only captures spatial dependencies across different geographic levels. The temporal information, which is also critical for spatio-temporal prediction tasks, is not represented except as a set of independent samples. Recent research has increasingly employed spatio-temporal transformers for spatio-temporal traffic forecasting, where a temporal transformer decoder follows the spatial transformer encoder to generate final predictions. These temporal attention layers capture the temporal dependencies in the input sequences. Temporal attention layers are typically preferred to conventional recurrent models due to their ability to learn long-distance dependencies and their efficient implementations in parallel settings. Unlike recurrent models, where information encoded at each time step is retained solely for the subsequent step and quickly diminishes in influence, attention layers calculate attention scores across the entire sequence concurrently, thereby promoting shared sequential information. However, GRU is still widely adopted in spatio-temporal prediction tasks (Lu etĀ al., 2022; Chen etĀ al., 2023; Ren etĀ al., 2023; Wang etĀ al., 2020). Additionally, our setting exhibits relatively short temporal lengths, and we find empirically that GRU outperforms a temporal transformer based model.
@@
Let be hours of input from geographic level . Then for each element in the input sequence, a GRU layer computes:
@@
where is the input at time , is the hidden state of the layer at time or the initial hidden state at time 0. The dimensions of the parameters are: . is the new hidden state at time , which is the hidden input for next time.
@@
In the GRU model, and changes the spatial dimension from coarse level to finer level . In our GRU model, we replace the three matrix multiplication with spatial attention layers. Then GRU is capable of modeling spatial and temporal dependencies concurrently. The calculation becomes:
where and are three separately trained spatial attention layers. We only use one GRU layer for our experiments.
@@ 4.4. Loss
We calculate loss at the target disaggregation level. However, since loss divides by the size of the vector, meaning that the loss will have different units at each aggregation level. For example, the loss at the TRACT level is in units of āevents per TRACT regionā while the loss at the BLOCK level is in units of āevents per BLOCK regionā, leaving the two losses incomparable. Therefore, to meaningfully compare the losses at different target disaggregation level, we normalize all loss terms to be events-per-region at a selected aggregation level. In our experiments, we normalize to the BLOCK level to assist with interpretation: we consider taxi rides per city block as more familiar than taxi rides per PUMA region, for example. We use the following loss at the target level :
@@ 5. Experiments
In this section, we describe the baseline models to compare against for the disaggregation task. Additionally, we describe the setup for the study of transfer learning and point synthesis.
@@ 5.1. Baseline Models
As our baseline, we test several heuristic-based disaggregation methods used commonly in the literature and in practice, as well as variants of more recent neural approaches:
ā¢
@@
ā¢
Convolutional Neural Network (CNN): The event count at each output region is extracted from the raster representation at the output aggregation level (e.g., city blocks). That raster representation in turn is learned from the raster representation of the input aggregation level (e.g., PUMA) using a typical UNet architecture(Ronneberger etĀ al., 2015). This approach models the non-linear relationships between regions, while emphasizing local effects through convolutional layers.
ā¢
@@
ā¢
Spatio-Temporal Attention (STA): a model with spatial and temporal attention layers. In contrast to typical spatio-temporal transformers employing multiple attention layers, our experiments employ a single spatial attention layer and a single temporal layer for two reasons: 1) The inclusion of a single spatial and temporal attention layer already results in a larger model compared to our proposed GRU (See Table 2). 2) Preliminary experiments indicate that adding more layers to STA detrimentally impacts performance.
We test three disaggregation tasks ā from PUMA to NTA, TRACT and BLOCK level respectively. We also tested GRU model without the containment maps as an ablation study. Batch size, learning rate, early stop tolerance are set to 32, 1e-4, and 50 respectively. We use 1xNVIDIA T4 GPU, 16 GB RAM, x4 vCPUs, for each task.
| Resolutions | FNN | GRU | STA | GRU |
| PUMA -Āæ NTA | 1.4K | 10K | 47K | 40K |
| PUMA -Āæ TRACT | 89K | 750K | 3.3M | 2.2M |
| PUMA -Āæ BLOCK | 15M | 128M | 585M | 377M |
@@ 5.2. Transfer Learning
As mentioned in §1, we do not always have access to sufficient historical data for training. In these cases, we can consider whether a model pre-trained on one variable can be used to make predictions for another variable. We can use our models in this transfer learning setting by training on, say, taxi data, then fine-tuning on a limited amount of bikeshare. This approach can significantly increase the value of open data, by allowing one dataset to be used in different applications. The idea is reasonable, given the interconnectedness of city dynamics as a complex system (Howe et al., 2022): everything depends on everything else. Moreover, the main signals being learned by any model are influenced by the built environment, human activity, and population density patterns, all of which tend to be exhibited in any spatio-temporal dataset in the same domain.
@@ 5.3. Event Data Synthesis
While the BLOCK aggregation level is high-resolution (smaller than census tracts for example), they remain spatially aggregated and cannot be directly utilized by applications requiring individual point data(Zhang etĀ al., 2020; Singh and Chen, 2023; Law etĀ al., 2018). Additionally, if we can synthesize individual taxi rides for NYC Taxi data after June 2016 given the aggregated taxi zone data, it would significantly enhance the utility of the datasets. A simple approach to synthesize points is to learn a high-resolution aggregation at BLOCK level, then randomly sample within these regions to generate individual points, considering the physical built-in environment (e.g., building, hotel etc.). If a region has events, we can sample events within its boundaries. Quality control is important for the synthesized data. Quantitatively, we can compare the distributions of real and disaggregated values across a set of BLOCK regions. We consider two statistics to measure similarities: Mutual Information (MI) and KL-divergence. We also assess the distribution similarity qualitatively through visualization.
@@ 6. Results
| Dataset | Resolutions | CW | AW | HR | CNN | FNN | GRU | STA | GRU | GRU |
| NYC Taxi | PUMA NTA | 1.8295 | 1.4072 | 0.4652 | 0.3601 | 0.2299 | 0.2208 | 0.2608 | 0.2125 | 0.2108* |
| PUMA TRACT | 1.8416 | 2.0651 | 0.9650 | 0.7297 | 0.5606 | 0.5492 | 0.6507 | 0.5565 | 0.5494 | |
| PUMA BLOCK | 3.0947 | 2.9705 | 1.7609 | 1.5306 | 1.4381 | 1.4130 | 1.4350 | 1.3904 | 1.3770* | |
| NYC Bikeshare | PUMA NTA | 0.3047 | 0.2071 | 0.1114 | 0.1012 | 0.0861 | 0.0843 | 0.0945 | 0.0843 | 0.0833* |
| PUMA TRACT | 0.5304 | 0.5032 | 0.2726 | 0.2595 | 0.2489 | 0.2427 | 0.2418 | 0.2374 | 0.2376* | |
| PUMA BLOCK | 1.4392 | 1.3792 | 0.3972 | 0.4065 | 0.3741 | 0.3668 | 0.3618 | 0.3599 | 0.3559* |
@@
Do neural methods improve performance over popular heuristic methods? (Yes) (Section §6.1)
Does our proposed GRU improve disaggregation performance relative to competing neural methods? (Yes) Does GRU converge faster than competing methods? (Yes) (Section §6.2)
@@ 6.1. All Neural Methods Outperform Heuristics
Table 3 presents the errors of the disaggregation task using both heuristic and neural methods. We observe that 1) As the target disaggregation resolution increases, the error increases across all methods, as expected. Predicting city block dynamics from large regional aggregates is a difficult problem. 2) Popular heuristic-based disaggregation methods, such as CW and AW, are non-competitive across all resolutions and datasets. The underlying heuristics are reasonable but cannot capture the dynamics at higher resolution: The distribution of taxi rides and bikeshare trips follows human activity patterns, which are influenced by the built environment, temporal effects, and non-linear interactions between regions, all of which are inaccessible to simple heuristics. 3) HR significantly outperforms other heuristics, indicating that temporal patterns are generally predictive. However, HR still overlooks the non-linear temporal and spatial variations present in the data, limiting its overall accuracy. 4) Neural-based methods outperform all commonly used heuristics in nearly all disaggregation tasks. GRU models, in particular, outperform models that solely rely on spatial information. GRU emerges as the best-performing baseline model. CNN models exhibit worse performance than HR models when disaggregating from PUMA all the way to BLOCK resolution. Upon investigation, this performance discrepancy is attributable to the information loss that occurs during the conversion from vector inputs to the image representations expected by CNN models. Overall, the results show that neural models, especially GRU models, outperform heuristic-based approaches.
6.2. GRU Improves Neural Results

### v4

@@ 3.1. Data Processing
We first define a 128256 grid and plot the regions at each geographic level. Each region has a corresponding value from the vector input — , and the number of pixels located within the region — . Then each pixel in each region takes the value of . Pixels shared by more than one region are assigned to whichever region is smaller to avoid ambiguity. The cost is that we lose information during translation since the raster representation does not represent region boundaries precisely, especially at BLOCK level.
We compute descriptive statistics, including means and standard deviations, for the counts of each dataset at different geographic levels. These statistics are shown in Table 1. For NYC Taxi data, the standard deviations are large demonstrating the significant variability in the count values between regions. For instance, Inwood in upper Manhattan has considerably fewer taxi rides compared to the more central areas of Manhattan. In the case of the NYC Bikeshare, we notice significantly smaller average values at the TRACT and BLOCK levels. This observation can be attributed to the distribution of the bikes, which may vary in frequency across different regions within the city. As a result, the average values of these datasets are lower at the more granular levels.
@@ 4. Models
In §4.1 we formally formulate the disaggregation problem. In §4.2, we introduce the structural property and containment maps that connect different levels of the aggregation hierarchy. In §4.3, we detail the architecture of the structurally-aware spatial attention (SASA) layer, including the calculation of global and structural attention using the containment maps. In §4.4, we present our proposed SARN model, which integrates SASA layers into the GRU layer. The model architecture is visualized in Figure 3
4.1. Notations & Problem Formulation
We adopt the following notation to be used throughout the paper:
@@ 4.1. Notations & Problem Formulation
The disaggregation problem is defined as — given a time series of traffic values in the past time steps at coarse level , we aim to learn a function that disaggregates the data to a finer level :
4.2. Structural Property — Containment
@@
The geographic levels we use exhibit structural problem containment — each region at level is wholly contained in a region at level such that regions are hierarchical. For example, one census tract comprises multiple census blocks. Assuming containment, we can identify a one-to-many map among the regions from any two geographic levels. Let be a binary matrix representing the containment relationship between geographic levels and . A value of 1 at position indicates that region at level is contained in region at level . Figure 4(a) shows an example where three regions from level at a single th timestamp are split into six regions from level . Figure 4(b) presents the corresponding binary containment map.
4.3. Structurally-Aware Spatial Attention Layer
@@
Spatio-temporal transformers have been applied to urban prediction tasks (Section 2). The multi-head spatial attention layers in the transformer blocks can capture complex spatial relationships among different regions. In our disaggregation task, we also leverage the attention mechanism to extract spatial information. The spatial attention is calculated across two geographic levels with different spatial dimensions. We propose a structurally-aware spatial attention (SASA) layer (see Figure 3(c)), which consists of multiple structurally-aware spatial attention heads (). Each head performs two attention calculations — a global attention and a structural attention. Given input from geographic level , the global attention score is calculated as:
The global attention enables each coarse region to attend to all finer regions, facilitating comprehensive interactions across different geographic levels. This macro-level interaction allows the model to capture broader dependencies. The structural attention (See Figure 4(c)) uses the containment map to force the attention to happen only among the connected regions:
The containment maps impose a constraint that requires the values of child regions to sum up to the value of their parent region. This ensures that the predicted values remain coherent and consistent across different geographic resolutions. By enforcing this hierarchical relationship, the model maintains the integrity of the spatial dependencies, allowing for accurate disaggregation results. The global and structural attention outputs are added and go through a feed-forward projection layer:
@@
A structurally-aware spatial attention layer (SASA) is composed with multiple spatial attention heads, enhancing the model capacity:
@@
The same calculation works with temporal inputs. Given hours of input , we reshape the dimension and calculate spatial attention based on Eq (7). In the regular attention calculation [44], a residual connection layer is applied after the multi-head attention layer: . In the disaggregation task, the input and output have different spatial dimensions. Therefore, residual connection is not applicable.
4.4. Structurally-Aware Recurrent Network
@@
The spatial attention layer only captures spatial dependencies across different geographic levels. The temporal information, which is also critical for spatio-temporal prediction tasks, is not represented except as a set of independent samples. Recent research has increasingly employed spatio-temporal transformers for spatio-temporal traffic forecasting. The temporal attention layers capture the temporal dependencies in the input sequences. Temporal attention layers are typically preferred to conventional recurrent models due to their ability to learn long-distance dependencies and their efficient implementations in parallel settings. Unlike recurrent models, where information encoded at each time step is retained solely for the subsequent step and quickly diminishes in influence, attention layers calculate attention scores across the entire sequence concurrently, thereby promoting shared sequential information. However, GRUs are still widely used in spatio-temporal prediction tasks (Lu et al., 2022; Chen et al., 2023; Ren et al., 2023; Wang et al., 2020). Additionally, our setting involves relatively short temporal lengths, and we empirically find that GRUs outperform temporal transformer-based models (see §6).
@@
The comparison of SARN and the original GRU model is visualized in Figure 3(a) and (b). Let be hours of input from geographic level . Then for each element in the input sequence, an original GRU layer computes:
@@
where is the input at time , is the hidden state of the layer at time or the initial hidden state at time 0. is the new hidden state, which is the hidden input for next time.
@@
In the GRU model, and changes the spatial dimension from coarse level to finer level . In our SARN model, we replace the three matrix multiplication with our SASA layers as introduced in §4.3. Then SARN is capable of modeling spatial and temporal dependencies concurrently (See Figure 3(b)). The calculation for the three gates becomes:
and are three separately trained spatial attention layers for the three gates.
@@ 4.5. Loss
We calculate loss at the target disaggregation level. However, since loss divides by the size of the vector, meaning that the loss will have different units at each aggregation level. For example, the loss at the TRACT level is in units of “events per TRACT region” while the loss at the BLOCK level is in units of “events per BLOCK region”, leaving the two losses incomparable. Therefore, to meaningfully compare the losses at different target disaggregation level, we normalize all loss terms to be events-per-region at a selected aggregation level. In our experiments, we normalize to the BLOCK level to assist with interpretation: we consider taxi rides per city block as more familiar than taxi rides per PUMA region, for example. We use the following loss at the target level :
@@ 5. Experiments
In this section, we describe the baseline models to compare against for the disaggregation task. Additionally, we describe the setup for the supplementary studies of downstream application, transfer learning and point synthesis.
@@ 5.1. Baseline Models
As our baseline, we test several heuristic-based disaggregation methods used commonly in the literature and in practice, as well as more advanced neural approaches:
@@
Convolutional Neural Network (CNN): The event count at each output region is extracted from the raster representation at the output aggregation level (e.g., city blocks). That raster representation in turn is learned from the raster representation of the input aggregation level (e.g., PUMA) using a typical UNet architecture(Ronneberger et al., 2015). This approach models the non-linear relationships between regions, while emphasizing local effects through convolutional layers.
@@
UrbanFM(Ouyang et al., 2020): a disaggregation model that generates fine traffic flow data from coarse, aggregated data using super-resolution technique. It adopts grid representations of traffic flow, which is similar to the CNN baseline. We modified the input PUMA image to be 444As mentioned in §2, their coarse observations are at pixel level, while our coarse values are at predifined geopolitical regions. When we process rasterized image inputs, pixels lying in the same region share the same value. Given this preprocessing, the normalization and distribution recovery layer can no longer be applied in the original UrbanFM model, which might negatively impact the results..
Spatio-Temporal Transformer (ST-Transformer): a model consisting a spatial and a temporal transformer. The spatial transformer shares the same spatial attention layer described in §4.3, with global attention only. The temporal transformer follows the set up in (Liu et al., 2023a) — given the output from the spatial transformer , we add hour-of-day and day-of-week index to the featurs, and generate a feature embedding. Then regular multi-head self-attention mechanism is applied. .
We test three disaggregation tasks — from PUMA to NTA, TRACT and BLOCK level respectively. Batch size, learning rate, early stop tolerance are set to 32, 1e-4, and 50 respectively. For STT model, the feature embedding size, number of temporal attention head, and number of temporal transformer layer are set to 64, 4, and 4 respectively. For our SARN, we only use one GRU layer, and each SASA layer uses 4 . We use 1xNVIDIA T4 GPU, 16 GB RAM, x4 vCPUs, for each task.
5.2. Downstream Application
To demonstrate the high quality and utility of the disaggregated data, we use traffic demand prediction as the downstream application task. Traffic demand prediction involves forecasting the number of vehicles required in a region within a specific time frame. Accurate traffic demand prediction enables efficient allocation of transportation resources, reducing mismatches between supply and demand, alleviating traffic congestion, and enhancing mobility service experiences. Coarse-level predictions (e.g., PUMA) are less useful as they tend to miss fine-grained human mobility patterns. Therefore, using disaggregated values at finer resolutions is more beneficial. The specific task is to predict the next hour’s taxi demand given the previous twelve hours of data. We first disaggregate taxi values from PUMA to NTA, TRACT, and BLOCK levels using our proposed SARN model on the original validation split. We then train several traffic prediction models (LSTM(Hochreiter and Schmidhuber, 1997), AGCRN(Bai et al., 2020), STSGCN(Song et al., 2020)) with both the disaggregated data and the original ground truth data at each level. Mean Absolute Error (MAE) is reported on the test set to evaluate performance.
@@ 5.3. Transfer Learning
As mentioned in §1, there are urban settings where the availability and quality of fine-grained data are limited for model training, posing a barrier to implement our model. This data availability issue is not just a limitation to our model, but across all neural models that require sufficient training; it is also universal to all urban computing tasks where neural models excel. In these cases, we can consider whether a model pre-trained on one variable can be used to make predictions for another variable. We can use our models in this transfer learning setting by training on, say, taxi data, then fine-tuning on a limited amount of bikeshare. This approach can significantly increase the value of open data, by allowing one dataset to be used in different applications. The idea is reasonable, given the interconnectedness of city dynamics as a complex system (Howe et al., 2022): everything depends on everything else. Moreover, the main signals being learned by any model are influenced by the built environment, human activity, and population density patterns, all of which tend to be exhibited in any spatio-temporal dataset in the same domain.
@@ 5.4. Event Data Synthesis
While the BLOCK aggregation level is high-resolution (smaller than census tracts for example), they remain spatially aggregated and cannot be directly utilized by applications requiring individual point data. Additionally, if we can synthesize individual taxi rides for NYC Taxi data after June 2016 given the aggregated taxi zone data, it would significantly enhance the utility of the datasets. A simple approach to synthesize points is to learn a high-resolution aggregation at BLOCK level, then randomly sample within these regions to generate individual points, considering the physical built-in environment (e.g., building, hotel etc.). If a region has events, we can sample events within its boundaries. Quality control is important for the synthesized data. Quantitatively, we can compare the distributions of real and disaggregated values across a set of BLOCK regions. We consider two statistics to measure similarities — Mutual Information (MI) and KL-divergence. MI measures the probabilistic dependency between two variables, where a higher value indicates greater shared information between the variables. KL divergence is a measure of entropy increase due to the use of an approximation distribution to the true distribution. A lower KL divergence value signifies a better match between the true distribution and our approximation. We also assess the distribution similarity qualitatively through visualization.
@@ 6. Results
Dataset | Resolutions | CW | AW | HR | UrbanFM | CNN | FNN | GRU | STT | SARN |
NYC Taxi | PUMA NTA | 1.8295 | 1.4072 | 0.4652 | 0.3989 | 0.3601 | 0.2300 | 0.2255 | 0.2729 |
0.2073* (
PUMA TRACT | 1.8416 | 2.0651 | 0.9650 | 0.8049 | 0.7324 | 0.5673 | 0.5631 | 0.6473 |
0.5443* (
PUMA BLOCK | 3.0947 | 2.9705 | 1.7609 | 1.8576 | 1.5306 | 1.4515 | 1.4278 | 1.4305 |
1.3763* (
NYC Bikeshare | PUMA NTA | 0.1678 | 0.1212 | 0.0619 | 0.0834 | 0.0587 | 0.0497 | 0.0478 | 0.0601 |
0.0473*(
PUMA TRACT | 0.2973 | 0.2945 | 0.1650 | 0.1787 | 0.1591 | 0.1497 | 0.1505 | 0.1584 |
0.1483*(
PUMA BLOCK | 0.7882 | 0.7574 | 0.2488 | 0.3726 | 0.2481 | 0.2492 | 0.2341 | 0.2513 |
0.2316*(
@@
Do neural methods improve performance over commonly used heuristic methods? (Yes) (Section §6.1)
Does our proposed SARN improve disaggregation performance relative to neural methods? (Yes) Does SARN converge faster than competing methods? (Yes) (Section §6.2)
Is the disaggragated data from SARN authentic and of great quality compared to real data? Can it be used in downstream applications? (Yes) (Section §6.3)
@@ 6.1. All Neural Methods Outperform Heuristics
Table 2 presents the errors of the disaggregation task using both heuristic and neural methods. We observe that:
As the target disaggregation resolution increases, the error increases across all methods, as expected. Predicting city block dynamics from large regional aggregates is a difficult problem.
Popular heuristic-based disaggregation methods, such as CW and AW, are non-competitive across all resolutions and datasets. While the underlying heuristics are reasonable, they cannot well capture the dynamics. The distribution of taxi rides and bikeshare trips follows human activity patterns influenced by temporal effects and non-linear interactions between regions, which simple heuristics cannot access.
HR significantly outperforms other heuristics (53.5% average increase over AW), indicating that temporal patterns are generally predictive. However, HR still overlooks the non-linear temporal and spatial variations present in the data, limiting its accuracy.
Neural-based methods outperform all common heuristics in nearly all disaggregation tasks. GRU models capture additional temporal dependencies, thus outperforming models that solely rely on spatial information. GRU emerges as the best-performing baseline model (2.3% average increase over FNN).
UrbanFM is outperformed by its counterpart image model CNN. We contribute the huge performance gap to the difference in setting of two disaggregation problems. The normalization and distribution recovery layer proposed for the original problem can no longer be applied in our setting, thus negatively impacting the results.
Overall, the results show that neural models, especially GRU models, outperform heuristic-based approaches.
6.2. SARN Improves Neural Results